OpenStudy (anonymous):

(-u^-4v^4) x -u^3v^4 / vu^0 someone please help me, I'm having a really hard time. I just need this one explained so I can do the rest on my own. Thank you

OpenStudy (anonymous):

first, is this your problem: \[\frac{ \left( -u^{-4}v^4 \right)\cdot-u^3v^4 }{ vu^0 }\text{?}\]

OpenStudy (anonymous):

yes it is